17. Find the solution of -4/3k-1/4≤5
Answer by RAY100(1637)

(-4/3)k -1/4 <=5
.
(-4/3)k <= 5.25
.
k>= 5.25(-3/4),,,,,,flip inequality when dividing by negative no.
.
k>= -3.9375,,,,(or -3 15/16)
.
check, let k=-3,,,,eqn is,,,(-4/3)(-3) -1/4 <= 5,,,4-1/4<=5,,,3.75<=5,,,ok